# DDPG-E2E: A Novel Policy Gradient Approach for End-to-End Communication Systems

### v1

@@ V-B Transmitter Training
Fig. 4 demonstrates the training of the transmitter (actor) and the critic networks, where $\bm{m}_{\mathrm{b}}$, $\bm{x}_{\mathrm{b}}$ and ${r}_{\mathrm{b}}$ are the sampled batches of messages, encoded signals, and rewards, respectively. The sampled batches are used to calculate the actor loss and the critic loss, which are then used for the networks’ optimizations. The actor calculates its loss by taking the negative expected value of the critic’s Q-value for the sampled message and encoded signal batches pair, while the critic calculates its loss by taking the mean squared error between its estimated Q-value for the sampled message and encoded signal batches pair. The actor network then updates its parameters using gradient ascent [32], while the critic network updates its parameters using gradient descent [33].
@@ V-C Receiver Training
Fig. 5 shows the training process of the receiver model, where $\bm{x}_{\mathrm{b}}$ denotes the sampled encoded signal batch from buffer, $\bm{y}_{\mathrm{b}}$ denotes the received signal batch after channel distortion, $\bm{m}_{\mathrm{b}}$ denotes the sampled message batch from buffer and $\hat{\bm{m}}_{\mathrm{b}}$ denotes the estimated message batch generated by the receiver. At every time step, the encoded signal batch $\bm{x}_{\mathrm{b}}$ and its corresponding message batch $\bm{m}_{\mathrm{b}}$ are randomly sampled from the experience buffer. The received signal batch $\bm{y}_{\mathrm{b}}$ and the sampled message batch $\bm{m}_{\mathrm{b}}$ are considered as the input-output pair for training the receiver model in a supervised manner. The estimated message batch $\hat{\bm{m}}_{\mathrm{b}}$ is then used to calculate the binary cross-entropy loss, given the sampled message batch $\bm{m}_{\mathrm{b}}$ as the label. The calculated loss is used to update the weights of the receiver using the Adam optimizer [31].
@@ VI Simulation Results
Figure 6: Episodic reward with (a) 128-bit and (b) 256-bit block length settings under Rayleigh and Rician fading channels.
In this section, we will evaluate the performance of our proposed DDPG-based solution under Rayleigh and Rician fading channel models. In particular, the Rayleigh fading channel describes the wireless communication scenarios where there is no direct path but only none-line-of-sight (NLOS) paths during signal propagation. The amplitude of the randomly scattering signals can be described using the probability density function (p.d.f.) of Rayleigh distribution, as follows:
@@
where $x$ is the amplitude of the signal, $\sigma$ is the scale parameter which determines the spread of Rayleigh distribution, $2{\sigma^{2}}=\mathbb{E}(x^{2})$ is the mean power of the signal and $\mathbb{E}(.)$ denotes statistical averaging. The Rician fading channel, on the other hand, describes the wireless communication scenario where there are both line-of-sight (LOS) and NLOS components during signal propagation. The p.d.f. of Rician distribution reflects the combined influence of LOS and NLOS effects on the signal amplitude, which can be expressed as:
@@
where $x$ is the amplitude of the signal, $\sigma$ is the scale parameter which determines the spread of Rician distribution, $L$ is the ratio of the power of the dominant path to the scattered path, $I_{0}$ is the modified Bessel function of the first kind.
The baselines for comparisons are the alternating training scheme proposed in [16] and the conditional GAN scheme proposed in [11]. The following simulations are evaluated by BER and BLER over different signal-to-noise ratio (SNR), ranging from 0 dB to 20 dB. The SNR is defined as:
$SNR=\frac{\mathbb{E}[\frac{1}{n}\left\lVert x\right\rVert^{2}_{2}]}{\sigma^{2}},$ (14)
where $\mathbb{E}[\frac{1}{n}\left\lVert x\right\rVert^{2}_{2}]$ is the expected energy per complex symbol, and $\sigma^{2}$ is the noise variance.
@@ VI-A Parameter Setting
For the hyper-parameters settings, the soft update rate $\tau$ is defined as 0.005 for slow updates of the target networks. The learning rates of the actor and critic networks are set as 0.0002 and 0.0001, respectively. The number of training episodes and time steps for every episode are set as 30,000 and 500, respectively. The discount factor is set at 0.1. The training of the DDPG-based solution is implemented on Rayleigh and Rician fading channels with different block sizes, i.e., 8, 128 and 256 information bits. The Rician factor is 1, indicating that the portions of the LOS components and NLOS components are equivalent. The simulation results with small block lengths are compared to that of the alternating training scheme, as this approach cannot be trained for data transmission with large block lengths. The results with large block lengths are compared to that of the conditional GAN solution since this approach can work with large block lengths.
@@ VI-B Performance Evaluation
Figure 7: BER vs. Iteration with block size of (a) 128 bits and (b) 256 bits.
Figure 8: BLER vs. SNR with block size of 8 and training SNR of (a) 20dB and (b) 10dB under Rayleigh fading channel.
Figure 9: BLER vs. SNR with block size of 8 and training SNR of (a) 20dB and (b) 10dB under Rician fading channel.
Figure 10: BLER vs. SNR with block size of 128 and training SNR of (a) 20dB and (b) 10dB under Rayleigh fading channel.
Figure 11: BLER vs. SNR with block size of 128 and training SNR of (a) 20dB and (b) 10dB under Rician fading channel.
Figure 12: BLER vs. SNR with block size of 256 and training SNR of (a) 20dB and (b) 10dB under Rayleigh fading channel.
Figure 13: BLER vs. SNR with block size of 256 and training SNR of (a) 20dB and (b) 10dB under Rician fading channel.
@@ VI-B1 Training Process
We first evaluate the training process of the proposed DDPG approach when training SNR of 20dB under 128-bit and 256-bit block length scenarios, as shown in Fig. 6(a) and Fig. 6(b), respectively. As can be seen from the figures, the average rewards in all cases can converge close to zero under both Rayleigh and Rician fading channels within 80 episodes. It is worth noting that the reward sent from the receiver to the transmitter is the negative value of the receiver’s training loss, as shown in (5). As such, an average reward approaching zero indicates a very high detection accuracy at the receiver. This verifies the feasibility of our proposed DDPG-based solution for training E2E communication systems without prior channel knowledge. In the following, we evaluate the detection performance of our proposed solution under different settings in terms of BLER and convergence rate.
@@ VI-B2 Convergence Rate
We conduct the convergence analysis by comparing the convergence rates of the proposed DDPG-based solution and the baseline schemes for data transmission with large block lengths (i.e., 128 bits and 256 bits) over Rayleigh fading and Rician fading channels, as shown in Fig. 7. Fig. 7(a) shows that the proposed DDPG-based scheme starts to converge to steady state after 3,000 iterations for both Rayleigh and Rician fading channels, while the baseline schemes cannot converge. Similarly, the proposed DDPG-based solution shows better convergence than the baseline schemes across the whole training process in the context of 256-bit data transmission. To complete the convergence analysis, we have measured the averaged running time of the proposed solution and the baseline schemes on Nvidia A100 Tensor Core GPU. The averaged running time per iteration of the proposed DDPG-based solution is about $2.687\times 10^{-2}$ seconds, while that of the GAN based approach and the RA-GAN based approach are about $4.490\times 10^{-2}$ seconds and $7.126\times 10^{-2}$ seconds, respectively. The result demonstrates the feasibility of the proposed DDPG-based scheme to quickly adapt to new channel scenario within a much shorter training time compared to the state-of-the-art solutions.
@@ VI-B3 Small Block Size
First, we consider the cases when transmitting 8 information bits per message and observe the performance of our proposed DDPG-based solution. The baseline schemes are the alternating training scheme proposed in [16], the conditional GAN based solution proposed in [11], and the residual-aided GAN based solution proposed in [13]. In particular, in Fig. 8, we vary the SNR and observe the BLER performance with train SNR of 10dB and 20dB. As can be seen from the figures, our proposed solution can obtain much better BLER performance compared to the all the baseline schemes. This demonstrates that the actor and critic networks in our proposed DDPG-based approach can efficiently learn the training process of the receiver and channel conditions through the feedback reward. Moreover, Fig. 8 suggests that training the proposed model with SNR = 10dB gives better performance than training with SNR = 20dB. This is because training the model at 20dB cannot cover low SNR ranges as good as training at 10dB which balances between low and high SNR ranges. In addition, training at 20dB can help the model to learn the characteristics of high SNR settings more efficiently than training at 10dB.
We then consider the Rician fading channel and observe the BLER performance of our proposed solution, as shown in Fig. 9. Similar to Rayleigh channel scenarios, our proposed approach can also achieve better BLER performance than all the baseline schemes. This demonstrates the effectiveness of our solution in different wireless settings. It is worth noting that the proposed DDPG algorithm is an online learning algorithm. As such, it can adjust its encoding policy at the transmitter very quickly (within 80 episodes as demonstrated in Fig. 6) based on the channel conditions.
@@ VI-B4 Large Block Size
Next, we compare our proposed solution with the GAN solution [^1] As mentioned, other approaches in the literature adopted the one-hot encoding mechanism that exponentially increases the size of input messages when increasing the block length. As such, for 128-bit and 256-bit messages, we only compare our solution with the conditional GAN approach and the residual-aided GAN approach. proposed in [11] and the RA-GAN solution proposed in [13] in terms of BLER with the block length of 128 and 256 information bits under the Rayleigh and Rician fading channels, as shown in Fig. 10, Fig. 11, Fig. 12, and Fig. 13, respectively. As can be observed, the proposed solution shows considerable improvement in BLER compared to the baseline schemes under both Rayleigh and Rician fading channels. The conditional GAN method utilizes conditional GANs to represent the channel effect in wireless communications, thereby connecting the transmitter and the receiver, which allows the gradient to be back-propagated from the receiver DNN to the transmitter DNN. However, the conditional GAN method cannot guarantee the availability of gradient when training the data with the large dimensional block length. The RA-GAN solution demonstrates slight performance improvement by utilizing ResNet and $\bm{l}_{2}$ regularization, but it cannot effectively mitigate the local optimization arising from BP. During the BP process from the receiver to the transmitter, via the generator-based channel, there is a certain degree of information loss in the error gradient, potentially leading to performance degradation. In contrast, the proposed DDPG-based approach circumvents the blocked gradient problem by using the reward feedback mechanism to jointly train the transmitter and the receiver networks. The reward signal feedback perfectly reserves the loss information of the receiver network, which explains the overall performance improvement of the proposed DDPG solution.
@@ VII Conclusion
In this article, we have proposed a DDPG-based E2E learning solution to relax the requirement of the prior channel model in conventional E2E communication systems. The proposed solution can also overcome the curse of dimensionality problem in E2E communication systems, which enables data transmission with large block length (i.e., 256 bits). In particular, with the DDPG algorithm, the transmitter can update its DNN by learning from the reward, i.e., loss value, sent from the receiver, given the current state, i.e., bitstream, and the chosen action, i.e., encoded symbol. In this way, implicit information about the training process of the receiver DNN can be learned by the transmitter to adapt its DNN, and thus improving the whole system’s performance. The simulation results in various wireless communication scenarios have demonstrated the effectiveness and significant improvement of our proposed solution compared to other solutions in the literature. In addition, the convergence analysis demonstrates better convergence of the proposed DDPG-based scheme compared to the baseline schemes across the whole training process.

### v2

@@ V-B Transmitter Training
Fig. 4 demonstrates the training of the transmitter (actor) and the critic networks, where $\bm{m}_{\mathrm{b}}$, $\bm{x}_{\mathrm{b}}$ and ${r}_{\mathrm{b}}$ are the sampled batches of messages, encoded signals, and rewards, respectively. The sampled batches are used to calculate the actor loss and the critic loss, which are then used for the networks’ optimizations. The actor calculates its loss by taking the negative expected value of the critic’s Q-value for the sampled message and encoded signal batches pair, while the critic calculates its loss by taking the mean squared error between its estimated Q-value and the target Q-value obtained from the Bellman function. The actor network then updates its parameters using gradient ascent [37], while the critic network updates its parameters using gradient descent [38].
@@ V-C Receiver Training
Fig. 5 shows the training process of the receiver model, where $\bm{x}_{\mathrm{b}}$ denotes the sampled encoded signal batch, $\bm{y}_{\mathrm{b}}$ denotes the received signal batch after channel distortion, $\bm{m}_{\mathrm{b}}$ denotes the sampled message batch and $\hat{\bm{m}}_{\mathrm{b}}$ denotes the estimated message batch generated by the receiver. At every time step, the encoded signal batch $\bm{x}_{\mathrm{b}}$ and its corresponding message batch $\bm{m}_{\mathrm{b}}$ are randomly sampled from the experience buffer. The received signal batch $\bm{y}_{\mathrm{b}}$ and the sampled message batch $\bm{m}_{\mathrm{b}}$ are considered as the input-output pair for training the receiver model in a supervised manner. The estimated message batch $\hat{\bm{m}}_{\mathrm{b}}$ is then used to calculate the binary cross-entropy loss, given the sampled message batch $\bm{m}_{\mathrm{b}}$ as the label. The calculated loss is used to update the weights of the receiver using the Adam optimizer [36].
@@ VI Simulation Results
Figure 6: (a) Averaged Episodic reward, and (b) Critic Loss Convergence with block length of 256 bits over fading channels.
In this section, we will evaluate the performance of our proposed DDPG-based solution under Rician, Ryaleigh, and 3GPP channel models. In particular, the Rayleigh fading channel describes the wireless communication scenarios where there is no direct path but only none-line-of-sight (NLOS) paths during signal propagation. The amplitude of the randomly scattering signals can be described using the probability density function (p.d.f.) of Rayleigh distribution, as follows:
@@
where $x$ is the amplitude of the signal, $\sigma$ is the scale parameter that determines the spread of Rayleigh distribution, $2{\sigma^{2}}=\mathbb{E}(x^{2})$ is the mean power of the signal, and $\mathbb{E}(.)$ denotes statistical averaging. The Rician fading channel, on the other hand, describes the wireless communication scenario where there are both line-of-sight (LOS) and NLOS components during signal propagation. The p.d.f. of Rician distribution reflects the combined influence of LOS and NLOS effects on the signal amplitude, which can be expressed as:
@@
where $x$ is the amplitude of the signal, $\sigma$ is the scale parameter that determines the spread of Rician distribution, $L$ is the ratio of the power of the dominant path to the scattered path, $I_{0}$ is the modified Bessel function of the first kind.
The baselines for comparisons are the alternating training (AT) scheme proposed in [16], the conditional GAN scheme proposed in [11] and the residual-aided GAN-based scheme proposed in [13]. The classic baseline considers LDPC channel coding for large block sizes, specifically 128 and 256 bits. Since LDPC requires a block size larger than 24 bits, it is only applied to these larger block sizes. The supervised learning baseline assumes that the channel gradients are known to the transmitter NN, with the transmitter and receiver together forming a complete autoencoder, referred to as AE. The learning rates of both the transmitter and the receiver are set as 0.001 for all the baselines. The following simulations are evaluated by BLER over different signal-to-noise ratios (SNRs), ranging from 0 dB to 20 dB.
@@ VI-A Parameter Setting
For the hyper-parameters settings, the soft update rate $\tau$ is set to 0.005 for slow updates of the target networks. The learning rates of the actor and critic networks are set as 0.002 and 0.001, respectively. Training consists of 30,000 episodes with 500 time steps per episode. The discount factor $\gamma$ is set at 0.01. The DDPG-based solution is trained on block fading channels with different block sizes, i.e., 8, 128, and 256 information bits. The Rician factor is 1, indicating that the portions of the LOS and NLOS components are equivalent.
@@ VI-B Performance Evaluation
Figure 7: (a) Receiver Loss Convergence for different schemes, and (b) Receiver Loss Convergence for different discount factors.
Figure 8: BLER vs. SNR with block size of 8 trained at (a) Rayleigh 20dB, (b) Rayleigh 10dB, (c) Rician 20dB, (d) Rician 10dB.
Figure 9: BLER vs. SNR with block size of 128 trained at (a) Rayleigh 20dB, (b) Rayleigh 10dB, (c) Rician 20dB, (d) Rician 10dB.
Figure 10: BLER vs. SNR with block size of 256 trained at (a) Rayleigh 20dB, (b) Rayleigh 10dB, (c) Rician 20dB, (d) Rician 10dB.
Figure 11: BLER vs. SNR on 3GPP.
@@ VI-B1 Training Process
We first evaluate the training process of the proposed DDPG approach trained at 20dB with 256-bit block length over both Rayleigh and Rician fading channels, as shown in Fig. 6. As can be seen in Fig. 6(a), the averaged episodic rewards can converge close to zero within 200 episodes. It is worth noting that the reward sent from the receiver to the transmitter is the negative value of the receiver’s training loss, as shown in (5). As such, an averaged episodic reward approaching zero indicates a high detection accuracy at the receiver. This verifies the effectiveness of our proposed DDPG-based solution for training E2E communication systems without prior channel knowledge. In Fig. 6(b), the critic loss converges within 700 episodes, indicating that the critic network effectively learns to minimize the difference between the target Q-value and the estimated critic value. This convergence enables the network to provide more accurate guidance to the transmitter, leading to better optimization of the system. In the following, we evaluate the detection performance of our proposed solution under different settings in terms of BLER and convergence rate.
@@ VI-B2 Convergence Rate
We conduct a convergence analysis by comparing the convergence rates of the receiver loss between the proposed DDPG-based solution and the baseline schemes, using a block size of 256 bits over a Rayleigh fading channel, as shown in Fig. 7(a). Clearly, the proposed DDPG-based scheme exhibits faster convergence compared to the baseline schemes. Specifically, the proposed DDPG scheme converges to approximately $3\times 10^{-5}$ for the receiver loss within 2,000 iterations, whereas the baseline methods do not achieve such low values. This demonstrates the effectiveness of the proposed DDPG-based scheme in rapidly adapting to new channel scenarios with significantly shorter training times compared to state-of-the-art solutions.
Next, we evaluate the loss convergence of the proposed DDPG-based solution across different discount factors $\gamma$, specifically evaluating values of 0.001, 0.01, 0.1, 0.5 and 0.9, as illustrated in Fig. 7(b). The results demonstrate that an increase in discount factor negatively impacts loss convergence. Notably, high discount factors, such as $\gamma=0.5$ and $\gamma=0.9$, exhibit significant divergence. In contrast, smaller discount factors, including 0.001, 0.01, and 0.1, can reach a similar steady state after 1,000 iterations. This behaviour suggests that a small discount factor better aligns with the multi-armed bandit scenario, where the adjacent transmitted signals are independent and uncorrelated.
@@ VI-B3 Small Block Size
First, we consider the cases when transmitting 8 information bits per message and observe the performance of our proposed DDPG-based solution. In particular, in Fig. 8(a), our proposed solution can obtain superior performance compared to other baselines. This demonstrates that the actor and critic networks in our proposed DDPG-based approach can efficiently learn the training process of the receiver and channel conditions through the feedback reward. Moreover, Fig. 8(b) suggests that training the proposed model at an SNR of 10 dB results in a flatter performance curve compared to training at 20 dB. This leads to better performance in the lower SNR range but comes at the cost of reduced performance in the high SNR regime.
We then consider the Rician fading channel and observe the BLER performance of our proposed solution, as shown in Fig. 8(c) and Fig. 8(d). Similar to Rayleigh channel scenarios, our proposed approach can also achieve better BLER performance than all the baseline schemes. This demonstrates the effectiveness of our solution in different wireless settings.
@@ VI-B4 Large Block Size
Next, we compare our proposed solution with the AT scheme [16], the GAN solution [11], the RA-GAN solution [13], and the LDPC classic baseline in terms of BLER with the block length of 128 and 256 information bits under the Rician and Rayleigh fading channels, as shown in Fig. 9 and Fig. 10, respectively. As can be observed, the proposed solution shows performance improvement in BLER compared to the baselines under both fading channels. The LDPC baseline performs better than RA-GAN and GAN schemes at low SNR, achieving lower BLER due to its optimization for noisy environments. However, RL-based schemes like AT and DDPG excel at high SNR levels, where they outperform LDPC. This is because DL-based schemes are trained at high SNR values, which may reduce their effectiveness at low SNR but enhance their performance at high SNR. LDPC codes, while effective in low SNR scenarios, become less efficient at high SNR due to increased complexity and decoding latency. As can be observed in Figs. 8, 9, and 10, the proposed DDPG solution can achieve a performance close to that of the supervised learning-based solution without requiring channel knowledge in advance.
@@
The conditional GAN method utilizes conditional GANs to represent the channel effect in wireless communications, thereby connecting the transmitter and the receiver, which allows the gradient to be back-propagated from the receiver NN to the transmitter NN. However, the conditional GAN method cannot guarantee the availability of gradient when training the data with the large dimensional block length. The RA-GAN solution demonstrates slight performance improvement by utilizing ResNet and $\bm{l}_{2}$ regularization, but it cannot effectively mitigate the local optimization arising from BP. During the BP process from the receiver to the transmitter, via the generator-based channel, there is a certain degree of information loss in the error gradient, potentially leading to performance degradation. In contrast, the RL-based schemes including AT scheme and the proposed DDPG-based approach circumvent the blocked gradient problem by using the loss feedback mechanism to jointly train the transmitter and the receiver networks. Specifically, the DDPG scheme adopts an actor-critic architecture to enable more effective training of the transmitter by leveraging the Q-value estimations generated by the critic network. The critic network is optimized by minimizing the difference between the reward batches and the predicted Q-values, where the reward batch accurately preserves the loss information from the receiver network. This mechanism explains the overall performance improvement of the proposed DDPG solution. Additionally, the experience replay buffer enhances learning stability and improves convergence by enabling the sampling of past experiences, with a focus on those associated with higher reward values.
VI-B5 Unknown Channel
Finally, we evaluate the performance of the proposed scheme under 3GPP channel setting [39]. In addition, we consider the case when the channel knowledge is not available in advance, which is common in practice where the channel between the transmitter and the receiver is typically complicated and may not be feasible to formulate. As shown in Fig. 11, the proposed DDPG-based solution demonstrates performance comparable to the classical AE baseline trained with a known channel. Notably, the performance advantage becomes obvious when compared to the AE baseline trained on an unknown channel. In specific, while the channel layer of the AE is fixed to a Rayleigh fading channel, the actual encoded signal passes through the 3GPP channel. The end-to-end loss at the receiver is used to update the entire AE via BP. This setup reflects the real-world scenario where the channel layer remains complex to be formulated. The results show that both the RL-based schemes and the LDPC scheme outperform AE trained on an unknown channel. In the low SNR range, the LDPC baseline shows the strongest resilience to noise, outperforming the other methods. However, as the SNR increases, both RL-based solutions surpass the LDPC scheme, demonstrating the effectiveness of the proposed CNN architecture and reinforcement learning procedure. The results underscore the ability of the proposed DDPG-based solution to adapt and excel under dynamic and complex channel scenarios.
@@ VII Conclusion
In this article, we have proposed a DDPG-based E2E learning solution to relax the requirement of the prior channel model in conventional E2E communication systems. In particular, the proposed CNN architecture can overcome the curse of dimensionality problem in E2E communication systems, which enables data transmission with large block lengths (i.e., 256 bits). With the DDPG algorithm, the transmitter updates its NN by learning from the reward (i.e., the loss value) sent by the receiver, based on the current state (i.e., bitstream) and the chosen action (i.e., encoded symbol). The use of a critic network for Q-value estimation, along with the experience replay buffer, enables the transmitter’s NN to more effectively learn the optimal mapping of encoded symbols. This process allows the transmitter to implicitly learn information about the receiver’s training, thereby adapting its NN and improving overall system performance. Simulation results across various wireless communication scenarios demonstrate the effectiveness of our proposed solution compared to existing methods. Furthermore, convergence analysis shows that the DDPG-based scheme achieves better convergence than baseline approaches throughout the entire training process.